# Remote State Estimation over Unreliable Channels with Unreliable Feedback: Fundamental Limits

### v1

@@ II-C A Team Decision-Making Problem
In our framework, the decision variables $\delta_{k}$ and $\hat{x}_{k}$ at each time $k$ for $k\in\mathbb{N}_{[0,N]}$ are determined by the encoder and the decoder in a distributed way, respectively. This distributed decision-making process relies on causal information. Let the information sets of the encoder and the decoder be denoted by $\mathcal{I}^{e}_{k}$ and $\mathcal{I}^{d}_{k}$. A coding policy profile $(\epsilon,\delta)$ consisting of a scheduling policy $\epsilon$ and an estimation policy $\delta$ is admissible if $\epsilon=\{\operatorname{\mathsf{P}}(\delta_{k}|\mathcal{I}^{e}_{k})\}_{k=0}^{N}$ and $\delta=\{\operatorname{\mathsf{P}}(\hat{x}_{k}|\mathcal{I}^{d}_{k})\}_{k=0}^{N}$, where $\operatorname{\mathsf{P}}(\delta_{k}|\mathcal{I}^{e}_{k})$ and $\operatorname{\mathsf{P}}(\hat{x}_{k}|\mathcal{I}^{d}_{k})$ are measurable stochastic kernels. In the sequel, we seek to find a globally optimal solution $(\epsilon^{\star},\delta^{\star})$ to a team decision-making problem cast as
$\underset{\epsilon\in\mathcal{E},\delta\in\mathcal{D}}{\operatorname{minimize}}\ \Phi(\epsilon,\delta)$ (11)
@@
$\Phi(\epsilon,\delta):=\operatorname{\mathsf{E}}\bigg{[}\sum_{k=0}^{N}\alpha_{k}\delta_{k}+(x_{k}-\hat{x}_{k})^{T}(x_{k}-\hat{x}_{k})\bigg{]}$ (12)
@@
The loss function in (12) is expressed in terms of two performance criteria: the mean-square-error and the packet-rate [^1] In the loss function in (12), the coefficient $1/N$ that appears in the definitions of these performance criteria is omitted, as it has no effect on the solutions.. The mean-square-error criterion, which is frequently used in the analysis of control systems, represents the quality of estimation. However, the packet-rate criterion, which is frequently used in the analysis of packet-switching networks, represents the cost of communication. We should emphasise that the problem (11) is a decision-making problem with a non-classical information structure. This problem is nonconvex and in general intractable. Finding a globally optimal solution $(\epsilon^{\star},\delta^{\star})$ to this problem allows us to determine the fundamental performance limits of the networked system in different communication regimes.
III Main Result
In this section, we present the main theoretical result on the synthesis of a globally optimal policy profile. Let $\check{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]$ and $\hat{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$. In addition, define the innovation at the encoder $\nu_{k}:=y_{k}-C_{k}\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k-1}]$, the estimation error at the decoder $\hat{e}_{k}:=x_{k}-\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$, the estimation mismatch $\tilde{e}_{k}:=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]-\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$, the conditional mean of the estimation mismatch at the encoder $\breve{e}_{k}:=\operatorname{\mathsf{E}}[\tilde{e}_{k}|\mathcal{I}^{e}_{k}]$, and the conditional covariance of the estimation mismatch at the encoder $R_{k}:=\operatorname{\mathsf{cov}}[\tilde{e}_{k}|\mathcal{I}^{e}_{k}]$.
In general, the information sets of the encoder and the decoder are specified as
$\mathcal{I}^{e}_{k}$ =$=\Big{\{}y_{t},z^{b}_{t},\delta_{s},\theta_{s}\big{|}t\in\mathbb{N}_{[0,k]},s\in\mathbb{N}_{[0,k-1]}\Big{\}}$
$\mathcal{I}^{d}_{k}$ =$=\Big{\{}z^{f}_{t},\delta_{s},\gamma_{s}\big{|}t\in\mathbb{N}_{[0,k]},s\in\mathbb{N}_{[0,k-1]}\Big{\}}$
The following definition introduces a value function associated with the loss function in (12).
@@ III Main Result
$V_{k}(\mathcal{I}^{e}_{k}):=$ $\min_{\epsilon\in\mathcal{E}:\delta=\delta^{\star}}\operatorname{\mathsf{E}}\bigg{[}\sum_{t=k}^{N-1}\alpha_{t}\delta_{t}+\hat{e}_{t+1}^{T}\hat{e}_{t+1}\Big{|}\mathcal{I}^{e}_{k}\bigg{]}$ (13)
@@
The value function $V_{k}(\mathcal{I}^{e}_{k})$ is a quantity defined from the perspective of the encoder, taking into account the fact that, due to causality, the scheduling policy $\epsilon\in\mathcal{E}$ at each time $k$ can affect $\hat{e}_{t}$ only from time $t=k+1$ onward.
@@
The tradeoff between the packet rate and the mean square error admits a globally optimal solution $(\epsilon^{\star},\delta^{\star})$ such that $\epsilon^{\star}$ is a symmetric threshold scheduling policy given by
$\delta_{k}=\mathds{1}_{\chi_{k}-\alpha_{k}\geq 0}$ (14)
in conjunction with $\check{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]$ for $k\in\mathbb{N}_{[1,N]}$, where $\chi_{k}(\breve{e}_{k},R_{k})=\lambda^{c}\breve{e}_{k}^{T}A_{k}^{T}A_{k}\breve{e}_{k}+\lambda^{c}\operatorname{tr}(A_{k}R_{k}A_{k}^{T})+\operatorname{\mathsf{E}}[V_{k+1}(\mathcal{I}^{e}_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}=0]-\operatorname{\mathsf{E}}[V_{k+1}(\mathcal{I}^{e}_{k+1})|\mathcal{I}^{e}_{k},\delta_{k}=1]$ is a symmetric function of $\breve{e}_{k}$ and a function $R_{k}$, which requires solving
$\check{x}_{k}$ $=A_{k-1}\check{x}_{k-1}+K_{k}\nu_{k}$ (15)
$Q_{k}$ $=\big{(}(A_{k-1}Q_{k-1}A_{k-1}^{T}+W_{k-1})^{-1}+L_{k}\big{)}^{-1}$ (16)
$\breve{e}_{k}$ $=(1-\bar{\theta}_{k-1})A_{k-1}\breve{e}_{k-1}+K_{k}\nu_{k}$ (17)
$R_{k}$ $=(1-\bar{\theta}_{k-1})A_{k-1}R_{k-1}A_{k-1}^{T}+\bar{\bar{\theta}}_{k-1}H_{k-1}$ (18)
for $k\in\mathbb{N}_{[1,N]}$ with initial conditions $\check{x}_{0}=m_{0}+K_{0}\nu_{0}$, $Q_{0}=(M_{0}^{-1}+C_{0}^{T}V_{0}^{-1}C_{0})^{-1}$, $\breve{e}_{0}=K_{0}\nu_{0}$, and $R_{0}=0$, where $K_{k}=Q_{k}C_{k}^{T}V_{k}^{-1}$, $L_{k}=C_{k}^{T}V_{k}^{-1}C_{k}$, $\bar{\theta}_{k-1}=\theta_{k-1}\delta_{k-1}\gamma_{k-1}+\lambda^{c}(1-\theta_{k-1})\delta_{k-1}$, $\bar{\bar{\theta}}_{k-1}=(1-\theta_{k-1})(\lambda^{c}\delta_{k-1}-{\lambda^{c}}^{2}\delta_{k-1})$, and $H_{k-1}=A_{k-1}\breve{e}_{k-1}\breve{e}_{k-1}^{T}A_{k-1}^{T}$; and $\delta^{\star}$ is a non-Gaussian linear estimation policy given by
$\hat{x}_{k}=A_{k-1}\hat{x}_{k-1}+\delta_{k-1}\gamma_{k-1}A_{k-1}(\check{x}_{k-1}-\hat{x}_{k-1})$ (19)
for $k\in\mathbb{N}_{[1,N]}$ with initial condition $\hat{x}_{0}=m_{0}$, where $\hat{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$, without being influenced by any signaling effect.
@@
See the Appendix. ∎

### v2

@@ II-C A Team Decision-Making Problem
In our framework, the decision variables $u_{k}$ and $\hat{x}_{k}$ at each time $k$ for $k\in\mathbb{N}_{[0,N]}$ are determined by the encoder and the decoder, respectively, as a team in a distributed way. This distributed decision-making process relies on causal information. Let the information sets of the encoder and the decoder be denoted by $\mathcal{I}^{e}_{k}$ and $\mathcal{I}^{d}_{k}$, which, in general, are specified as
$\mathcal{I}^{e}_{k}$ $=\Big\{y_{t},z^{b}_{t},u_{s},\theta_{s}\big|t\in\mathbb{N}_{[0,k]},s\in\mathbb{N}_{[0,k-1]}\Big\},$ (11)
$\mathcal{I}^{d}_{k}$ $=\Big\{z^{f}_{t},u_{s},\gamma_{s}\big|t\in\mathbb{N}_{[0,k]},s\in\mathbb{N}_{[0,k-1]}\Big\}.$ (12)
A coding policy profile $(\epsilon,\delta)$ consisting of a scheduling policy $\epsilon$ and an estimation policy $\delta$ is admissible if $\epsilon=\{\operatorname{\mathsf{P}}(u_{k}|\mathcal{I}^{e}_{k})\}_{k=0}^{N}$ and $\delta=\{\operatorname{\mathsf{P}}(\hat{x}_{k}|\mathcal{I}^{d}_{k})\}_{k=0}^{N}$, where $\operatorname{\mathsf{P}}(u_{k}|\mathcal{I}^{e}_{k})$ and $\operatorname{\mathsf{P}}(\hat{x}_{k}|\mathcal{I}^{d}_{k})$ are measurable stochastic kernels. In the sequel, we seek to find a globally optimal solution $(\epsilon^{\star},\delta^{\star})$ to a team decision-making problem cast as
$\underset{\epsilon\in\mathcal{E},\delta\in\mathcal{D}}{\operatorname{minimize}}\ \Phi(\epsilon,\delta),$ (13)
@@
$\Phi(\epsilon,\delta):=\operatorname{\mathsf{E}}\bigg[\sum_{k=0}^{N}\alpha_{k}u_{k}+(x_{k}-\hat{x}_{k})^{T}(x_{k}-\hat{x}_{k})\bigg],$ (14)
@@
The loss function in (14) is expressed in terms of two performance criteria: mean square error and packet rate. Note that we omitted the coefficient $\frac{1}{N+1}$ in these criteria, as it does not affect the solutions to the problem in (13). The mean-square-error criterion, which is frequently used in the analysis of control systems, represents the quality of estimation. However, the packet-rate criterion, which is frequently used in the analysis of packet-switching networks, represents the cost of communication. Note that, apart from the time averaging, the expectation is adopted in the definition of the loss function in order to formulate an optimisation problem whose solution is not path-dependent (see e.g., [70]). We should also emphasise that the problem in (13) is a decision-making problem with a non-classical information structure. This problem is nonconvex and in general intractable. Finding a globally optimal solution $(\epsilon^{\star},\delta^{\star})$ to this problem enables us to determine the fundamental performance limits of the networked system.
The next definition introduces a value function related to the loss function in (14), which will be instrumental in our analysis.
@@
$V_{k}(\mathcal{I}^{e}_{k}):=$ $\min_{\epsilon\in\mathcal{E}:\delta=\delta^{\star}}\operatorname{\mathsf{E}}\bigg[\sum_{t=k}^{N-1}\alpha_{t}u_{t}+\hat{e}_{t+1}^{T}\hat{e}_{t+1}\Big|\mathcal{I}^{e}_{k}\bigg],$ (15)
@@
Remark 5
The value function $V_{k}(\mathcal{I}^{e}_{k})$, defined over the space of the encoder’s information set $\mathcal{I}^{e}_{k}$, represents the minimum cumulative loss incurred from time $k$ to the terminal time $N$, under the assumption that the estimation policy $\delta^{\star}$ is fixed and given by $\hat{x}_{k}=\operatorname{\mathsf{E}}[x_{k}\mid\mathcal{I}^{d}_{k}]$. This value function is a quantity from the perspective of the encoder, and incorporates the impact of the optimal scheduling decisions. Note that, due to causality, the scheduling policy $\epsilon\in\mathcal{E}$ at each time $k$ can affect $\hat{e}_{t}$ only from time $t=k+1$ onward. This implies that at the final time step $k=N$, the encoder has no further opportunity to influence the estimation process through its scheduling decision. Hence, $u_{N}$ should be equal to zero in the optimal scheduling policy $\epsilon^{\star}$.
III Main Results
In this section, we present the main theoretical results on the synthesis of a globally optimal policy profile. Let $\check{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]$ and $\hat{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$, and define the innovation from the perspective of the encoder $\nu_{k}:=y_{k}-C_{k}\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k-1}]$, the estimation error from the perspective of the decoder $\hat{e}_{k}:=x_{k}-\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$, the estimation mismatch $\tilde{e}_{k}:=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]-\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$, the conditional mean of the estimation mismatch from the perspective of the encoder $\breve{e}_{k}:=\operatorname{\mathsf{E}}[\tilde{e}_{k}|\mathcal{I}^{e}_{k}]$, the conditional covariance of the estimation mismatch from the perspective of the encoder $R_{k}:=\operatorname{\mathsf{cov}}[\tilde{e}_{k}|\mathcal{I}^{e}_{k}]$, and the value function residual $\Delta_{k}:=\operatorname{\mathsf{E}}[V_{k+1}(\mathcal{I}^{e}_{k+1})|\mathcal{I}^{e}_{k},u_{k}=0]-\operatorname{\mathsf{E}}[V_{k+1}(\mathcal{I}^{e}_{k+1})|\mathcal{I}^{e}_{k},u_{k}=1]$.
@@ III Main Results
The causal tradeoff between the packet rate and the mean square error under unreliable communication with unreliable acknowledgements admits a globally optimal solution $(\epsilon^{\star},\delta^{\star})$ such that $\epsilon^{\star}$ is a symmetric threshold-based scheduling policy given by
$u_{k}=\mathds{1}_{\chi_{k}-\alpha_{k}\geq 0},$ (16)
for $k\in\mathbb{N}_{[0,N-1]}$, in conjunction with $\check{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{e}_{k}]$, where $\chi_{k}(\breve{e}_{k},R_{k})=\lambda^{c}\breve{e}_{k}^{T}A_{k}^{T}A_{k}\breve{e}_{k}+\lambda^{c}\operatorname{tr}(A_{k}R_{k}A_{k}^{T})+\Delta_{k}$ is a symmetric function of $\breve{e}_{k}$ and a function of $R_{k}$, which requires solving
$\check{x}_{k}$ $=A_{k-1}\check{x}_{k-1}+K_{k}\nu_{k},$ (17)
$Q_{k}$ $=\big((A_{k-1}Q_{k-1}A_{k-1}^{T}+W_{k-1})^{-1}+L_{k}\big)^{-1},$ (18)
$\breve{e}_{k}$ $=(1-\bar{\theta}_{k-1})A_{k-1}\breve{e}_{k-1}+K_{k}\nu_{k},$ (19)
$R_{k}$ $=(1-\bar{\theta}_{k-1})A_{k-1}R_{k-1}A_{k-1}^{T}+\bar{\bar{\theta}}_{k-1}H_{k-1},$ (20)
for $k\in\mathbb{N}_{[1,N]}$ with initial conditions $\check{x}_{0}=m_{0}+K_{0}\nu_{0}$, $Q_{0}=(M_{0}^{-1}+C_{0}^{T}V_{0}^{-1}C_{0})^{-1}$, $\breve{e}_{0}=K_{0}\nu_{0}$, and $R_{0}=0$, where $\nu_{k}=y_{k}-C_{k}A_{k-1}\check{x}_{k-1}$, $K_{k}=Q_{k}C_{k}^{T}V_{k}^{-1}$, $L_{k}=C_{k}^{T}V_{k}^{-1}C_{k}$, $\bar{\theta}_{k-1}=\theta_{k-1}u_{k-1}\gamma_{k-1}+\lambda^{c}(1-\theta_{k-1})u_{k-1}$, $\bar{\bar{\theta}}_{k-1}=(\lambda^{c}-{\lambda^{c}}^{2})(1-\theta_{k-1})u_{k-1}$, and $H_{k-1}=A_{k-1}\breve{e}_{k-1}\breve{e}_{k-1}^{T}A_{k-1}^{T}$; and $\delta^{\star}$ is a linear switching estimation policy given by
$\hat{x}_{k}=A_{k-1}\hat{x}_{k-1}+u_{k-1}\gamma_{k-1}A_{k-1}\tilde{e}_{k-1},$ (21)
for $k\in\mathbb{N}_{[1,N]}$ with initial condition $\hat{x}_{0}=m_{0}$, where $\hat{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$, without being influenced by implicit information.
@@
The proof proceeds as follows. We first establish that, without loss of optimality, the encoder can transmit the conditional mean $\check{x}_{k}=\mathbb{E}[x_{k}\,|\,\mathcal{I}^{e}_{k}]$ and the decoder can compute the state estimate $\hat{x}_{k}=\mathbb{E}[x_{k}\,|\,\mathcal{I}^{d}_{k}]$. We then complete the argument in three steps:
  1. (i)
    In the first step, we show that $\Phi(\epsilon^{n},\delta^{o})=\Phi(\epsilon^{o},\delta^{o})$, where $\epsilon^{n}$ is an innovation-based scheduling policy, i.e., it can be expressed in the form of $\operatorname{\mathsf{p}}(u_{k}|\boldsymbol{\nu}_{0:k},\boldsymbol{z}^{b}_{0:k},\boldsymbol{u}_{0:k-1},\boldsymbol{\theta}_{0:k-1})$.
  2. (ii)
    In the second step, we show that $\Phi(\epsilon^{s},\delta^{o})\leq\Phi(\epsilon^{n},\delta^{o})$, where $\epsilon^{s}$ is a special form of $\epsilon^{n}$ that is symmetric with respect to $\boldsymbol{\nu}_{0:k}$.
  3. (iii)
    In the final step, we show that $\Phi(\epsilon^{\star},\delta^{\star})\leq\Phi(\epsilon^{s},\delta^{o})$, where $\epsilon^{\star}$ is a special form of $\epsilon^{s}$ that satisfies the Bellman optimality equation and $\delta^{\star}$ is the same as $\delta^{o}$, constructed based on $\hat{x}_{k}=\operatorname{\mathsf{E}}[x_{k}|\mathcal{I}^{d}_{k}]$.
See the Appendix for the complete proof. ∎